2fn-f(n-2) = fn+1 for n>3


1
Expert's answer
2021-08-23T17:48:00-0400

The recursive definition for generating Fibonacci numbers and the Fibonacci sequence is:


"f_n = f_{n-1} + f_{n-2} \ where\ n\geq3"

Then


"f_{n+1} = f_{n} + f_{n-1}"

"f_{n-1}=f_{n+1}-f_{n}"

Substitute


"f_n = f_{n-1} + f_{n-2}=f_{n+1}-f_{n}+ f_{n-2}"

Therefore


"2f_n-f_{n-2} =f_{n+1}, n\geq3,\ True"